Required information [The following information applies to the questions displayed below.) On January 1, 2021, Splash City is can someone solve this for me? thks
Answer #1

Par value of bonds = $480,000

Semi annual interest payment = 480,000 x 7% x 6/12

= $16,800

Effective interest rate = 8%

Semi annual Effective interest rate = 4%

Issue price of bonds = $447,385

Discount on bonds payable = Par value of bonds - Issue price of bonds

= 480,000 - 447,385

= $32,615

Date Cash paid Interest expense Change in carrying value Carrying value
1/1/21 447,385
6/30/21 16,800 447,385 x 4% = 17,895 17,895 - 16,800 = 1,095 447,385 + 1,095 = 448,480
12/31/21 16,800 448,480 x 4% = 17,939 17,939 - 16,800 = 1,139 448,480 + 1,139 = 449,619

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubt. Thanks.